OpenStudy (anonymous):

Solve the proportion: 4/21 = x/168

OpenStudy (anonymous):

\[\frac{4}{21}\text{=}\frac{x}{168}\]\[21x=168*4 \]\[x=\frac{168*4}{21}=32\]